OpenStudy (anonymous):

Given that the tangent line to y=f(x) at (4,3) passes through the point (0,2) , find f(4) and f'(4).

OpenStudy (dumbcow):

its given that f(4) = 3 f'(4) is the slope of the line that goes through the 2 points m = (3-2)/(4-0) = 1/4 f'(4) = 1/4